In a multiple-choice test, each question has four options. students will get 4 points for each correct answer; lose 1 point for each incorrect answer; and receive no points for unanswered questions. A student does not know the correct answer for one question. Is it to her advantage or disadvantage to guess an answer? Show your calculations for expected value. Show your work for full credit.
Please help!!!!!!!!!!!!!!!! I am lost, you guys are the best here and I have learned a lot but stuck on this one.

Answer by stanbon(57290)

In a multiple-choice test, each question has four options. students will get 4 points for each correct answer; lose 1 point for each incorrect answer; and receive no points for unanswered questions.
---------------------------------------------------------------
A student does not know the correct answer for one question. Is it to her advantage or disadvantage to guess an answer? Show your calculations for expected value. Show your work for full credit.
-----------
E(x) = 4(1/4) + (-1)(3/4) + 0(0) = 1-(3/4) = 1/4 point.
--------------------
The student should guess.
---------------------
Cheers,
Stan H.